OpenStudy (anonymous):

how many degrees are in the measure of an interior angle of a regular 7 sided polygon

OpenStudy (anonymous):

The measure of any exterior angle of a regular \(n\)-gon is \(\dfrac{360^\circ}{n}\). Thus the measure of each interior angle, which is supplementary to any given exterior angle, is \(\left(180-\dfrac{360}{n}\right)^\circ\).

OpenStudy (anonymous):

Thank you so much ..I'm new to this and it has become very helpful (:

OpenStudy (anonymous):

You're welcome